OpenStudy (anonymous):

complete the square... 9x^2 - 12x = 14

OpenStudy (anonymous):

Complete Solution: \[ 9x^2 - 12x - 14 = 0 \rightarrow x^2 - \frac {4}{3}x - \frac {14}{9} = 0 \rightarrow \left( x - \frac {2}{3} \right)^2 - \dfrac {10}{9} = 0 \rightarrow \left( x - \frac {2}{3} \right)^2 = \dfrac {10}{9} \rightarrow \cdots \]The rest is straight-forward. Don't forgot to take the positive and negative square roots! :)